# Cover times with stochastic resetting

### v1

@@ 3.2. Proofs
Under the assumptions of section 3, let $\Omega\subset M\backslash X_{r}(0)$ and let $p^{*}=p^{*}(r)$ be any function of $r>0$ satisfying $p^{*}\sim p_{\Omega}$ as $r\to\infty$. Then for integers $m\geq 1$, we have
$\mathbb{E}[(\tau_{r}(\Omega))^{m}]\sim\frac{m!}{(rp^{*})^{m}}\quad\text{as }r\to\infty.$
@@
We start with the lower bound: Let $x^{*}\in U_{T}^{\text{far}}$. The survival probability of $T_{r}$ is that of the union of hitting times to each point in the target region, which itself has a natural lower bound,
(14) $\mathbb{P}(T_{r}>t)=\mathbb{P}(\cup_{x\in U_{T}}\{\tau_{r}(B(x,R))>t\})\geq\mathbb{P}(\tau_{r}(B(x^{*},R))>t).$
@@
$\geq\int_{0}^{\infty}\mathbb{P}(\tau_{r}(B(x^{*},R))>z^{1/m})\,\textup{d}z=\mathbb{E}[(\tau_{r}(B(x^{*},R)))^{m}].$
@@
$\liminf_{r\to\infty}\mathbb{E}[(rp_{B(x^{*},R)}T_{r})^{m}]\geq\lim_{r\to\infty}\mathbb{E}[(rp_{B(x^{*},R)}\tau_{r}(B(x^{*},R)))^{m}]=m!.$
Hence, for $p^{*}\sim p_{B(x^{*},R)}$ as $r\to\infty$, we have
$\liminf_{r\to\infty}\mathbb{E}[(rp^{*}T_{r})^{m}]\geq m!.$
Determining the limit supremum is more involved: Let $\varepsilon\in(0,1)$ and cover the compact target region $U_{T}\subseteq M$ using a finite set $\{x_{1},\dots,x_{k}\}\in M\backslash U_{T}^{\text{far}}$ such that the assumption in (13) is satisfied with $\Omega_{0}=U_{T}^{\text{far}}$ and $\Omega_{1}=\{x_{1},\dots,x_{k}\}$ and further that
@@
$=\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\mathbb{E}[(\tau_{r}(\cup_{i\in I}B(x_{i},(1-\varepsilon)R)))^{m}]\Big{)}$
Multiplying through by $(rp_{B(x^{*},(1-\varepsilon)R)})^{m}$ yields
(17) $\begin{split}\mathbb{E}[(r&p_{B(x^{*},(1-\varepsilon)R)}T_{r})^{m}]\\ &\leq\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\mathbb{E}[(rp_{B(x^{*},(1-\varepsilon)R)}\tau_{r}(\cup_{i\in I}B(x_{i},(1-\varepsilon)R)))^{m}]\Big{)}\\ &=\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\Big{(}\frac{p_{B(x^{*},(1-\varepsilon)R)}}{p_{\cup_{i\in I}B(x_{i},(1-\varepsilon)R)}}\Big{)}^{m}\\ &\qquad\qquad\qquad\qquad\times\mathbb{E}[(rp_{\cup_{i\in I}B(x_{i},(1-\varepsilon)R)}\tau_{r}(\cup_{i\in I}B(x_{i},(1-\varepsilon)R)))^{m}]\Big{)}\end{split}$
Using Lemma 2 and the assumption in (13) to take the limit as $r\to\infty$ yields
(18) $\limsup_{r\to\infty}\mathbb{E}[(r$ $p_{B(x^{*},(1-\varepsilon)R)}T_{r})^{m}]\leq 0.$
Since the left-hand side of (18) is trivially non-negative, the proof is complete.
@@ 4.1. Definitions and main result
Let $X_{r}$ be the process $X$ modified to instantaneously reset to its initial position $i_{0}\in\Gamma$ according to $\sigma$ as in (12). Assume the detection ‘radius’ of $X_{r}$ is zero; that is, the searcher sees only the node it occupies. Then the region covered by the resetting searcher up through time $t>0$ is
$S_{r}(t):=\cup_{s=0}^{t}X_{r}(s)\subseteq\Gamma,$
and the corresponding cover time of the target region $U_{T}\subseteq\Gamma$ is
@@
$p_{\Omega}:=\mathbb{P}(\tau_{0}(\Omega)<\sigma).$
@@
Let $d(\Gamma_{0},\Gamma_{1})\in\mathbb{N}$ denote the length of the geodesic path(s) from a set of nodes $\Gamma_{0}\subset\Gamma$ to a set of nodes $\Gamma_{1}\subset\Gamma$,
$d(\Gamma_{0},\Gamma_{1}):=\min\{n:\mathcal{P}\in\Gamma^{n+1},\mathcal{P}(1)\in\Gamma_{0},\mathcal{P}(n+1)\in\Gamma_{1}\}.$
We define the set of all paths from $\Gamma_{0}$ to $\Gamma_{1}$ with this minimum length $d(\Gamma_{0},\Gamma_{1})$,
$\mathcal{S}(\Gamma_{0},\Gamma_{1}):=\{\mathcal{P}\in\Gamma^{n+1}:\mathcal{P}(1)\in\Gamma_{0},\mathcal{P}(n+1)\in\Gamma_{1},n=d(\Gamma_{0},\Gamma_{1})\}.$
@@
Finally, let $d^{*}\geq 1$ denote the length of the geodesic path(s) between $i_{0}$ and the furthest node(s) in the finite target set $U_{T}$,
$d^{*}:=\sup_{j\in U_{T}}d(i_{0},j),$
@@
$U_{T}^{\text{far}}:=\{j\in U_{T}:d(i_{0},j)=d^{*}\}\subseteq U_{T}.$
Unlike in previous sections, we do not require that there be a unique furthest target point in $U_{T}$ from the initial and resetting position, or equivalently that $|U_{T}^{\text{far}}|=1$.
@@
$\mathbb{E}[(T_{r})^{m}]\sim\frac{m!\,K_{m}}{r^{m(1-d^{*})}}\quad\text{as }r\to\infty$
@@
$\mathbb{E}[(T_{r})^{m}]\sim\frac{m!}{(rp_{0})^{m}}\quad\text{as }r\to\infty,$
@@
$p_{0}=\frac{\Lambda(U_{T}^{\text{far}})}{r^{d^{*}}}\sim p_{U_{T}^{\text{far}}}\quad\text{as }r\to\infty.$
Further, $rp_{0}T_{r}$ converges in distribution to a unit rate exponential random variable,
$rp_{0}T_{r}\to_{\textup{dist}}\textup{Exponential}(1)\quad\text{as }r\to\infty.$
4.2. Proofs
@@
(22) $\begin{split}\mathbb{P}(T_{r}>t)&=\mathbb{P}(\cup_{j\in U_{T}}\{\tau_{r}(j)>t\})\\ &=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\mathbb{P}(\cap_{j\in J}\{\tau_{r}(j)>t\})\Big{)}\\ &=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\mathbb{P}(\tau_{r}(J)>t)\Big{)}\end{split}$
@@
$\mathbb{E}[(T_{r})^{m}]=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\mathbb{E}[(\tau_{r}(J))^{m}]\Big{)}.$
For any set of nodes $J\in\Gamma$, it was recently shown that $p_{J}\sim\Lambda(J)/r^{d(i_{0},J)}$ as $r\to\infty$ [15]. This result combined with Lemma 2 and the assumption that $|U_{T}|<\infty$ yields
$\lim_{r\to\infty}\mathbb{E}[r^{m(1-d^{*})}T_{r}^{m}]$ $=\lim_{r\to\infty}\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\frac{r^{-md^{*}}}{p_{J}^{m}}\mathbb{E}[(rp_{J}\tau_{r}(J))^{m}]\Big{)}$
$=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\lim_{r\to\infty}\frac{r^{-md^{*}}}{p_{J}^{m}}\lim_{r\to\infty}\mathbb{E}[(rp_{J}\tau_{r}(J))^{m}]\Big{)}$
$=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\lim_{r\to\infty}\frac{r^{m(d(i_{0},J)-d^{*})}}{(\Lambda(J))^{m}}$

### v2

@@ 2.2. Proof
Let $p_{\Omega}$ be as in (3) and let $\Omega\subset M$ be any union of balls such that $X_{r}(0)\not\in\Omega$. Then for integers $m\geq 1$, we have
$\mathbb{E}[(\tau_{r}(\Omega))^{m}]\sim\frac{m!}{(rp_{\Omega})^{m}}\quad\text{as }r\to\infty.$
@@
We start with the lower bound: Let $y\in U_{T}$. The survival probability of $T_{r}$ is that of the union of hitting times to each point in the target region, which itself has a natural lower bound,
(8) $\mathbb{P}(T_{r}>t)=\mathbb{P}(\cup_{x\in U_{T}}\{\tau_{r}(B(x,R))>t\})\geq\mathbb{P}(\tau_{r}(B(y,R))>t).$
@@
$\geq\int_{0}^{\infty}\mathbb{P}(\tau_{r}(B(y,R))>z^{1/m})\,\textup{d}z=\mathbb{E}[(\tau_{r}(B(y,R)))^{m}].$
@@
$\liminf_{r\to\infty}\mathbb{E}[(rp_{B(y,R)}T_{r})^{m}]\geq\lim_{r\to\infty}\mathbb{E}[(rp_{B(y,R)}\tau_{r}(B(y,R)))^{m}]=m!.$
Hence
$\liminf_{r\to\infty}\mathbb{E}[(rp_{B(y,R)}T_{r})^{m}]\geq m!.$
Determining the limit supremum is more involved: Let $\varepsilon\in(0,1)$ and $x^{*}\in U_{T}^{\textup{far}}$, and cover $U_{T}$ by a finite set of $\varepsilon R$-radius balls with centers $\{x_{1},\dots,x_{k}\}\in U_{T}$,
@@
$=\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\mathbb{E}[(\tau_{r}(\cup_{i\in I}B(x_{i},(1-\varepsilon)R)))^{m}]\Big{)}.$
Multiplying through by $(rp_{B(x^{*},(1-2\varepsilon)R)})^{m}$ yields
(11) $\begin{split}\mathbb{E}[(r&p_{B(x^{*},(1-2\varepsilon)R)}T_{r})^{m}]\\ &\leq\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\mathbb{E}[(rp_{B(x^{*},(1-2\varepsilon)R)}\tau_{r}(\cup_{i\in I}B(x_{i},(1-\varepsilon)R)))^{m}]\Big{)}\\ &=\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\Big{(}\frac{p_{B(x^{*},(1-2\varepsilon)R)}}{p_{\cup_{i\in I}B(x_{i},(1-\varepsilon)R)}}\Big{)}^{m}\\ &\qquad\qquad\qquad\qquad\times\mathbb{E}[(rp_{\cup_{i\in I}B(x_{i},(1-2\varepsilon)R)}\tau_{r}(\cup_{i\in I}B(x_{i},(1-\varepsilon)R)))^{m}]\Big{)}.\end{split}$
Using Lemma 2 and the assumption (4) to take the limit as $r\to\infty$ yields
(12) $\limsup_{r\to\infty}\mathbb{E}[(r$ $p_{B(x^{*},(1-2\varepsilon)R)}T_{r})^{m}]\leq 0.$
Since the left-hand side of (12) is trivially non-negative, the proof is complete. ∎
@@ 3.1. Definitions and main result
Let $X_{r}$ be the process $X$ modified to instantaneously reset to its initial position $i_{0}\in\Gamma$ according to $\sigma$ as in (1). Assume the detection ‘radius’ of $X_{r}$ is zero; that is, the searcher detects only the node it occupies. Then the region covered by the resetting searcher up through time $t>0$ is
$S_{r}(t):=\cup_{s=0}^{t}X_{r}(s)\subseteq\Gamma$
and the corresponding cover time of the finite target region $U_{T}\subseteq\Gamma$ is
@@
(13) $p_{\Omega}:=\mathbb{P}(\tau_{0}(\Omega)<\sigma).$
@@
Let $h(\Gamma_{0},\Gamma_{1})\in\mathbb{N}$ denote the length of the geodesic path(s) from a set of nodes $\Gamma_{0}\subset\Gamma$ to a set of nodes $\Gamma_{1}\subset\Gamma$,
$h(\Gamma_{0},\Gamma_{1}):=\min\{n:\mathcal{P}\in\Gamma^{n+1},\mathcal{P}(1)\in\Gamma_{0},\mathcal{P}(n+1)\in\Gamma_{1}\}.$
We define the set of all paths from $\Gamma_{0}$ to $\Gamma_{1}$ with this minimum length $h(\Gamma_{0},\Gamma_{1})$,
(16) $\mathcal{S}(\Gamma_{0},\Gamma_{1}):=\{\mathcal{P}\in\Gamma^{n+1}:\mathcal{P}(1)\in\Gamma_{0},\mathcal{P}(n+1)\in\Gamma_{1},n=h(\Gamma_{0},\Gamma_{1})\}.$
@@
Finally, let $h^{*}\geq 1$ denote the length of the geodesic path(s) between $i_{0}$ and the furthest node(s) in the finite target set $U_{T}$,
(18) $h^{*}:=\sup_{j\in U_{T}}h(i_{0},j)$
@@
(19) $U_{T}^{\text{far}}:=\{j\in U_{T}:h(i_{0},j)=h^{*}\}\subseteq U_{T}.$
@@
$\mathbb{E}[(T_{r})^{m}]\sim\frac{m!\,K_{m}}{r^{m(1-h^{*})}}\quad\text{as }r\to\infty$
@@
$\mathbb{E}[(T_{r})^{m}]\sim\frac{m!}{(rp^{*})^{m}}\quad\text{as }r\to\infty$
@@
$p^{*}=\frac{\Lambda(U_{T}^{\text{far}})}{r^{h^{*}}}\sim p_{U_{T}^{\text{far}}}\quad\text{as }r\to\infty.$
Further, $rp^{*}T_{r}$ converges in distribution to a unit rate exponential random variable,
$rp^{*}T_{r}\to_{\textup{dist}}\textup{Exponential}(1)\quad\text{as }r\to\infty.$
Note that using (7) on Theorem 3 to estimate the optimal resetting rate on a network yields the approximation $r_{\textup{opt}}\approx 0$ despite stochastic resetting having the capacity to reduce the MCT. Below we provide an alternative estimate of the optimal resetting rate on a network as well as a criterion for when inducing minimal stochastic resetting into the search process reduces the MCT. In particular, the proof of Theorem 3 reveals that the MCT can be described exactly by MFPTs to target nodes,
(20) $\mathbb{E}[T_{r}]=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\mathbb{E}[\tau_{r}(J)]\Big{)}.$
Supplementing a recent derivation [32] with a higher order term, we find that MFPTs under infrequent exponentially-distributed resetting satisfy
(21) $\mathbb{E}[\tau_{r}(J)]=\mathbb{E}[\tau_{0}(J)]+\frac{r}{2}\Phi(J)+\frac{r^{2}}{2}\Big{(}\frac{1}{3}\mathbb{E}[\tau_{0}^{3}(J)]+\Phi(J)\mathbb{E}[\tau_{0}(J)]\Big{)}+o(r^{2})$
where
$\Phi(J):=(\mathbb{E}[\tau_{0}(J)])^{2}-\text{Var}(\tau_{0}(J))$
and $\textup{Var}(\tau)$ denotes the variance of $\tau$. Hence by substitution of (21) into (20),
(22) $\begin{split}\mathbb{E}[T_{r}]&=\mathbb{E}[T_{0}]+\frac{r}{2}\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\Phi(J)\Big{)}\\ &\quad+\frac{r^{2}}{2}\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\big{(}\frac{1}{3}\mathbb{E}[\tau_{0}^{3}(J)]+\Phi(J)\mathbb{E}[\tau_{0}(J)]\big{)}\Big{)}+o(r^{2}).\end{split}$
Infrequent resetting thus reduces the MCT on a network if
(23) $\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\Phi(J)\Big{)}<0.$
If the target is merely one node (i.e. $|U_{T}|=1$), we recover the corresponding criterion for reducing the MFPT in [32] which requires that the coefficient of variation of the first passage time without resetting be greater than one,
$\textup{CV}(\tau_{0}(U_{T})):=\frac{\sqrt{\textup{Var}(\tau_{0}(U_{T}))}}{\mathbb{E}[\tau_{0}(U_{T})]}>1.$
Assuming the criterion in (23) is satisfied, we can use the zero of the $r$-derivative of (22) to attain an approximation of the optimal resetting rate to minimize MCTs on a network,
$r_{\textup{opt}}\approx\frac{\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j}\sum_{J\subseteq U_{T},\,|J|=j}\Phi(J)\Big{)}}{2\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\big{(}(1/3)\mathbb{E}[\tau_{0}^{3}(J)]+\Phi(J)\mathbb{E}[\tau_{0}(J)]\big{)}\Big{)}}.$
3.2. Proof
The proof of Theorem 3 relies on the following lemma which was proven in [28] (see Proposition 1 in [28]). Lemma 5 itself uses that the distribution of the search time without resetting evolves according to
(24) $\mathbb{P}(\tau_{0}(\Omega)\leq t)\sim\frac{\Lambda(\Omega)}{h(i_{0},\Omega)!}t^{h(i_{0},\Omega)}\quad\text{as }t\to 0^{+}$
as shown in Proposition 1 of [31].
Lemma 5.
Let $\Omega\subset\Gamma$ and assume (24). Then, for integers $m\geq 1$,
$\mathbb{E}[(\tau_{r}(\Omega))^{m}]\sim\frac{m!}{(rp_{\Omega})^{m}}\quad\text{as }r\to\infty$
and $p_{\Omega}$ in (13) satisfies
$p_{\Omega}\sim\frac{\Lambda(\Omega)}{r^{h(i_{0},\Omega)}}\quad\text{as }r\to\infty.$
@@ 3.2. Proof
(25) $\begin{split}\mathbb{P}(T_{r}>t)&=\mathbb{P}(\cup_{j\in U_{T}}\{\tau_{r}(j)>t\})\\ &=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\mathbb{P}(\cap_{j\in J}\{\tau_{r}(j)>t\})\Big{)}\\ &=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\mathbb{P}(\tau_{r}(J)>t)\Big{)}.\end{split}$
@@
(26) $\mathbb{E}[(T_{r})^{m}]=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\mathbb{E}[(\tau_{r}(J))^{m}]\Big{)}.$
Lemma 5 and the assumption that $|U_{T}|<\infty$ yields
$\lim_{r\to\infty}\mathbb{E}[r^{m(1-h^{*})}T_{r}^{m}]$ $=\lim_{r\to\infty}\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\frac{r^{-mh^{*}}}{p_{J}^{m}}\mathbb{E}[(rp_{J}\tau_{r}(J))^{m}]\Big{)}$
$=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\lim_{r\to\infty}\frac{r^{-mh^{*}}}{p_{J}^{m}}\lim_{r\to\infty}\mathbb{E}[(rp_{J}\tau_{r}(J))^{m}]\Big{)}$
$=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\lim_{r\to\infty}\frac{r^{m(h(i_{0},J)-h^{*})}}{(\Lambda(J))^{m}}$